Version 1 (modified by mggr, 15 years ago) (diff)

partially complete, paused to fix broken networking

Meeting notes: 06 May 2006

Present: Sue, Dom, Fabio, Siva, Mike, Peter.

Agenda items


Status of MOLES and CSML population


  • MOLES: still waiting on internal catalogue population
  • CSML: 1/3 of COAPEC done (ticket:48), will probably do sample sets for ERA40 (ticket:49) and ACSOE (ticket:50). Dome noted data quality issues.


  • Helen is doing the population and wasn't present to report. Fabio believes it's still to be done. NOCS is also still waiting on OCCAM data from BODC.


  • Siva nearly done on alpha CSML tickets (ticket:51 & ticket:52).
  • CSML manually generated, just needs to create Stub_B's and DIFs.
  • Some delays with eXist due to changing guest pw and effects of this on the XQueries.


  • Produced level3 sample CSML for AVHRR, will do the other sensors shortly (ticket:53)
  • Will make sure we have matching MOLES for this CSML by Alpha.

Other metadata issues

  • Sue checking if the new GCMD namespace is suitable to recomment for use by DPs (and later provide guidelines), ticket:227
  • MOLES->DIF transformation (ticket:281)
    • Siva currently using a jar Kev gave him that converts directly from MOLES to DIF
    • Sue to check with Kev and Roy about whether we can produce Stub-Bs, and whether we still need to (XSL update)
  • Examples to be provided for security in MOLES/CSML records


  • PML to ensure they have matching MOLES records for their sample CSML.
  • DPs to update metadata population tickets with progress.
  • Sue to test GCMD namespace (ticket:227)
  • Sue checking with Kev about ticket:281 (Stub-B XSL update)
  • Dom adding a ticket regarding CSML updates and examples including security (done, ticket:319)